OpenStudy (anonymous):

help.. i need to find a fuction rule for table (0,7) (1,6) (2,5) (3,4) ??

OpenStudy (anonymous):

hmmm it looks like the x and y values add up to 7 [BIG HINT]

OpenStudy (anonymous):

can you write the rule now?

OpenStudy (anonymous):

for the first two but it dosent work for the last two??

OpenStudy (anonymous):

what rule u using ?

OpenStudy (anonymous):

thats the problem i dont even know how to write the rule i like completly forgot how to do these, can you give me a example maybe

OpenStudy (anonymous):

ok.. the question says a function rule. so i think it wants f(x) = some rule

OpenStudy (anonymous):

now, remember the hint i gave at the beginning? how are the x and y coordinates related?

OpenStudy (anonymous):

soo f(x)=7 ? but 7 dosent apply to the last 2 coordinates

OpenStudy (anonymous):

how would you write this out mathematically: "x and y coordinates add up to 7"

OpenStudy (anonymous):

if you can answer this question ^^^ you're almost there...

OpenStudy (anonymous):

yeah i dont think im almost there haha, well x+7=y?? or f(7)=....

OpenStudy (anonymous):

no... that's not how to translate "x AND y coordinates add up to 7"

OpenStudy (anonymous):

shouldn't the 7 be on the right hand side? and the x and y on the left joined by "+" ???

OpenStudy (anonymous):

hmm x+y=7 ?

OpenStudy (anonymous):

yes!!!

OpenStudy (anonymous):

now, solve for y... by subtracting x from both sides... what u got now?

OpenStudy (anonymous):

well for the first one i got 0=0 the 6=6 is tht right so far

OpenStudy (anonymous):

no... with the equation you gave: x + y = 7, i want you to solve for y... (get the y by itself on one side of the equal sign)

OpenStudy (anonymous):

y=7-x or y=-x+7

OpenStudy (anonymous):

great!!! there's your rule: f(x) = 7 - x or f(x) = -x + 7

OpenStudy (anonymous):

it works for all those points....

OpenStudy (anonymous):

yeah im pretty sure, too check i just plug in the x coordinates right

OpenStudy (anonymous):

correct... i'm glad u didn't give up... you did great! :)

OpenStudy (anonymous):

ahh thanks for the help!!! i should have knew this im in algebra 2 i just forget sometimes

OpenStudy (anonymous):

happens to everyone.... and yw....:)